### v2

@@ A.2.2 Experiment Setup
Following recent worksOreshkin et al. (2019); Nie et al. (2022); Zhang & Yan (2023); Zeng et al. (2023); Zhou et al. (2021); Liu et al. (2024), we conducted our experiments on four subsets of ETT as well as the Electricity dataset. The input length was fixed to 96 as the previous work did and the prediction lengths were between {96,192,336,720} respectively. Note that a randomized sparse version of confident pairwise regression learning was used when the prediction length was 720 for ETT and 192/336/720 for Electricity. We adopted two state-of-the-art time series models, including iTransformerLiu et al. (2024) and MinusformerLiang et al. (2024) as strong baselines to show the generalizability of our proposed AdaPRL method. The pairwise loss type was set to MAE to reduce penalty over outliers, and a coarse-grained grid search was performed on a weight factor $\alpha$ between {0.5,1.0,2.0,4.0,8.0,16.0,32.0,64.0}.
@@ A.2.3 Overall Performance
As was shown in Table 8 and Table 9, by comparing using single MSE loss, our proposed method, AdaPRL, substantially achieved higher precision in every subset of ETT datasets ETTm1, ETTm2, ETTh1, ETTh2 across all prediction length. Two experiment results also proved that AdaPRL could work well with different model backbone, including two state-of-the-art time-series forecasting models, iTransformer and Minusformer, and boosted the baseline to even higher performance. The improvement of MSE and MAE of AdaPRL in the entire ETT dataset was 1.65% and 1.86% for iTransformer, and 1.19% and 1.52% for Minusformer. The improvement of MSE and MAE of AdaPRL in the ECL dataset was 1.73% and 2.33% for iTransformer, and 0.74% and 1.80% for Minusformer respectively. Visualization examples of both datasets nad models were illustrated in section B.7.
A.3 Age Prediction datasets
@@ A.3.1 Datasets
The AgeDB  [^5] https://ibug.doc.ic.ac.uk/resources/agedb/ Moschoglou et al. (2017) is a dataset released by Sberbank AI Lab, focused on advancing research in cross-age facial recognition. It consists of 16,488 facial images, each annotated with gender and age. We use the RetinaFace model for face detection and cropping to ensure each image fully captures the face. The images are resized to 256x256 pixels and the pixel values are normalized. The dataset is split into training, testing, and validation sets in a 6:2:2 ratio.
The CLAP2016  [^6] https://chalearnlap.cvc.uab.cat/dataset/26/description/ Escalera et al. (2016) dataset is a large-scale facial image database designed for the estimation of both real and apparent age. It contains 7,591 facial images, each annotated with the corresponding real and apparent age. According to the official data format, these images are divided into 4,113 training images, 1,500 validation images, and 1,978 test images. Similarly, we employ the RetinaFace model for facial image detection and cropping, and apply the same preprocessing techniques as those used in AgeDB for subsequent image processing.
@@ A.3.3 Overall Performance
The detailed experimental results were delineated in Table 10. We highlighted the highest-performing datasets in boldface and the second-best in underlined text. Unlike other regression tasks that may concurrently evaluate MSE and MAE metrics, age estimation typically prioritized the MAE metric. We augmented L1, Huber, and L2 loss functions with both original pairwise learning (L2+PRL) and AdaPRL to assess performance across two datasets. The experimental findings yielded the following insights: within the AgeDB and CLAP2016 datasets, AdaPRL consistently attained the most exemplary performance irrespective of the baseline regression loss employed. Specifically, on the AgeDB dataset, the MAE metric was reduced by 1.13%, 0.36%, and 0.71% when combined with L1, Huber, and L2 regressions, respectively; on the CLAP2016 dataset, reductions were 7.47%, 6.13%, and 4.13%, respectively. Additionally, while achieving more accurate age estimations, both PRL and AdaPRL also garnered improvements in the Kendall’s $\tau$ across both datasets. This demonstrated that our approach excelled in preserving the ordinal relationships of actual ages, thereby more precisely reflecting the relative age dynamics among individuals.
@@
Table 10: Comparison of various methods’ performance on the AgeDB and CLAP2016 datasets. Rel.Imp denoted the relative improvement over the corresponding metrics.
A.4 Natural language understanding datasets
@@ A.4.2 Experiment Setup
In our experiments, we employ pre-trained RoBERTa Liu et al. (2019) base model as the baseline backbone. To better facilitate training in larger batch sizes and improve accuracy, we use the LAMB You et al. (2020) optimizer with a batch size of 320 and a learning rate of 8e-4 using a warm-up cosine annealing scheduler. We implement an early stopping strategy based with patience of 2 in the validation set. A grid search was conducted on the hyperparameters including ranking loss type, $\alpha$ between {0.001,0.002, 0.005, 0.01,0.02,0.05,0.1,0.2} and $\theta$ between {0,0.3,0.9}. The experiments are carried out over 10 runs to reduce randomness and the average performance on the metrics of the test set is reported.
@@ A.4.3 Overall Performance
The detailed experimental results were presented in Table 11. Similarly, we emphasized the top-performing datasets in bold and denoted the second-best performing datasets with underlines. The findings demonstrated that the proposed AdaPRL(L2) loss function excelled, achieved the lowest MSE of 0.3561 and MAE of 0.2922, represented improvements of 3.39% and 2.92% over the standard L2 loss, respectively. Although the original pairwise learning (L2+PRL) did not match AdaPRL’s performance, it still offered a significant enhancement over the L2 loss in both MSE and MAE. Additionally, we observed that both AdaPRL and orginal pairwise learning (L2+PRL) attained a higher Kendall’s $\tau$ (0.6858) compared to L2, further illustrated that our approach more effectively maintained the relative ranking among samples when predicted the toxicity levels of user comments, which was crucial for accurately assessed the severity of reviews.
@@
Table 11: Regression accuracies for Pairwise Regression Learning and AdaPRL on Jigsaw. Rel.Imp indicated relative improvement over corresponding metrics. The experiments were repeated 10 times, and the average results were reported.
@@ A.5.1 Datasets
Janestreet  [^8] https://www.kaggle.com/competitions/jane-street-real-time-market-data-forecasting/dataDesai et al. (2024) data comprise a set of time series with 79 anonymized features, 1 sample weight column and 9 responders, representing real market data. The goal of is to forecast one of the responders for the future. We formalize the task as a univariate regression problem, skip the first 500 days of data to avoid missing data, and randomly split 50% / 50% of the data as validation and test set for the last 100 days. Following the data description and the official evaluation protocol, we use weighted $R^{2}$ as evaluation metrics.
@@ A.5.2 Experiment Setup
In our experiments, we employed a multi-head attention modelVaswani et al. (2023) as encoder followed by a multilayer perceptron as the baseline backbone. The input sequence length, stride, and prediction length were set to 64, 16 and 1 respectively, and the model was trained on respond_6 as a single target. All experiments used AdamWLoshchilov (2017) as optimizer and the weighted MSE as $L_{reg}$. We implemented an early stopping strategy based with patience of 3 in the validation set. A grid search was conducted on the hyperparameters including ranking loss type, $\alpha$ between {0.005, 0.01,0.02,0.04,0.08,0.16,0.32}. The experiments were conducted over 3 runs and we reported the average weighted $R^{2}$ metrics on the test set.
@@ A.5.3 Overall Performance
The comprehensive experimental results were delineated in Table 12. The premier method AdaPRL attained the highest weighted $R^{2}$ score of 0.00617, thereby eclipsing alternative approaches. In comparison to the baseline loss function weighted MSE, which yielded a weighted $R^{2}$ of 0.00540, AdaPRL demonstrated a relative enhancement of 14.81%. Furthermore, the pairwise learning technique weighted MSE+PRL achieved a marginally superior performance of 0.00575; however, it did not match the precision of AdaPRL. These findings substantiated that AdaPRL effectively increased regression accuracy, surpassing both the conventional weighted MSE and the original pairwise learning-based approaches.
@@
Table 12: Regression accuracies for Pairwise Regression Learning (PRL) and AdaPRL on Janestreet. Rel.Imp indicated relative improvement over corresponding metrics. The experiments were repeated three times, and the average results were reported.
@@ A.6.2 Overall Performance
As shown in Table 13, AdaPRL achieved the highest accuracy among various loss functions on our industrial dataset, demonstrating the effectiveness of the proposed training techniques. Native pairwise regression learning improved MSE and MAE by 0.51% and 1.60%, respectively. In contrast, the AdaPRL method, leveraging pairwise learning and uncertainty estimation, achieved total gains of 0.96% and 1.93% in MSE and MAE. Notably, in large-scale music streaming applications with tens of millions of users, even a modest accuracy improvement of 0.1% is empirically significant and can positively impact business outcomes. The proposed method showcased substantial superiority in capturing and utilizing the relative relationships within user data. Furthermore, Kendall’s $\tau$ score for the proposed approach was the highest among the evaluated loss functions, reflecting enhanced ranking capabilities, which are crucial for identifying high-value users in real-world applications.
@@
Table 13: Regression accuracies for Pairwise Regression Learning (PRL) and AdaPRL on large-scale proprietary industrial regression dataset from a major music streaming app. Rel.Imp indicated relative improvement over corresponding metrics. The experiments were repeated three times and the average results were reported.
@@ B.1 Sensitivity of hyperparameter alpha
To investigate the impact and sensitivity of the hyperparameter in AdaPRL, we conducted experiments on the hyperparameter alpha in this section in the MovieLens-1M, KuaiRec and ETT datasets. We employed an RMSE type AdaPRL as the auxiliary loss for the MovieLens-1M dataset and MAE type AdaPRL for all the other datasets. $\theta$ were set to 0 for all experiments. All other experimental settings remained consistent with those described above.
@@
The results were shown in Figures 2 and 3. As was illustrated in the figures, in most cases of alpha, AdaPRL consistently achieved higher precision than the regression baseline across the MovieLens-1M, KuaiRec, and ETT datasets. These results suggested that our proposed approach was robust and not highly sensitive to the choice of the alpha hyperparameter.
@@ B.2 Effect of sparsity in AdaPRL
As we had analyzed before in 2.3.2, if randomness was induced for $L_{SCPRL}$ when generating the sparse area $S$, this method could be viewed as a variant of dropout applied to confidence pairwise learning. Therefore, we found it essential to conduct a quantitative analysis of the potential effects of dropout on the algorithm and empirically investigated how the underlying mechanisms of random dropout influenced the model. The evaluation was conducted on a subset of the ETT dataset, specifically focusing on prediction tasks with a forecasting horizon of 720 time steps in this experiment to analysis since the problem of bottleneck of memory and computation was mainly raised in the scenario of multivariate time series forecasting while incorporating with AdaPRL. We used iTransformer as baseline model and conducted a grid search of sparsity based on random masking techniques, hyperparameter alpha was optimized via a coarse-grained grid search and the best result was reported. All other settings remained the same as was described above in A.2.2.
@@
The result was shown in Figure 4. It was obvious that as the sparsity of area gradually decreased from 0.9 to 0.1, where computation gradually increased, the algorithm’s evaluation loss MAE and MSE exhibited an overall downward trend. This indicated that reducing sparsity enhanced the model’s predictive accuracy and robustness, likely by preserving essential information and mitigating the loss of critical features. However, when the sparsity was set to 1.0, corresponding to using a non-sparse algorithm, both MAE and MSE were worse than those observed at most other sparsity levels. This suggested that the introduction of a certain (probably relatively lower) level of sparsity, corresponding to small drop rate in dropout, played a crucial role in preventing overfitting and reducing computational complexity, thereby improving generalization performance. Conversely, the absence of sparsity, although it computed the relative difference of more pairs of data and required more memory consumption, still led to a suboptimal performance compared with a sparse version of AdaPRL. The full result of sparsity test of ETT dataset was listed in Table 14.

### v3

@@ A.2.2 Experiment Setup
Following recent works Oreshkin et al. (2019); Nie et al. (2022); Zhang & Yan (2023); Zeng et al. (2023); Zhou et al. (2021); Liu et al. (2024), we conducted our experiments on four subsets of ETT as well as the Electricity dataset. The input length was fixed to 96 as the previous work did, and the prediction lengths were between {96,192,336,720} respectively. Note that a randomized sparse version of confident pairwise regression learning was used when the prediction length was 720 for ETT and 192/336/720 for Electricity. We adopted two state-of-the-art time series models, including iTransformer Liu et al. (2024) and Minusformer Liang et al. (2024) as strong baselines to show the generalizability of our proposed AdaPRL method. The pairwise loss type was set to MAE to reduce penalty over outliers, and a coarse-grained grid search was performed on a weight factor $\alpha$ between {0.5,1.0,2.0,4.0,8.0,16.0,32.0,64.0}.
@@ A.2.3 Overall Performance
As shown in Table 6, Table 7, Table 8 and Table 9, comparing with using a single MSE loss, our proposed method AdaPRL significantly improved the precision across all prediction lengths for every subset of the ETT dataset (ETTm1, ETTm2, ETTh1, ETTh2) as well as the ECL dataset. The experimental results also demonstrated that AdaPRL worked well with different model backbones, including two state-of-the-art time series forecasting models, iTransformer and Minusformer, and boosted their performance to even higher levels. AdaPRL improved the MSE and MAE by 1.65% and 1.86% with iTransformer, and 1.19% and 1.52% with Minusformer on the entire ETT dataset. On the ECL dataset, AdaPRL improved the MSE and MAE by 1.73% and 2.33% with iTransformer, and 0.74% and 1.80% with Minusformer, respectively. The visualizations of the forecasting results and uncertainty prediction for these datasets and models are provided in section B.7.
A.3 Age Prediction
@@ A.3.1 Datasets
The AgeDB [^5] https://ibug.doc.ic.ac.uk/resources/agedb/ Moschoglou et al. (2017) is a dataset released by Sberbank AI Lab, focused on advancing research in cross-age facial recognition. It consists of 16,488 facial images, each annotated with gender and age. We use the RetinaFace model for face detection and cropping to ensure each image fully captures the face. The images are resized to 256x256 pixels, and the pixel values are normalized. The dataset is split into training, testing, and validation sets in a 6:2:2 ratio.
The CLAP2016666https://chalearnlap.cvc.uab.cat/dataset/26/description/ Escalera et al. (2016) dataset is a large-scale facial image database designed for the estimation of both real and apparent age. It contains 7,591 facial images, each annotated with the corresponding real and apparent age. According to the official data format, these images are divided into 4,113 training images, 1,500 validation images, and 1,978 test images. Similarly, we employ the RetinaFace model for facial image detection and cropping, and apply the same preprocessing techniques as those used in AgeDB for subsequent image processing.
@@ A.3.3 Overall Performance
The detailed experimental results are delineated in Table 10. We highlighted the highest-performing datasets in bold and the second-best in underlined text. Unlike other regression tasks that can concurrently evaluate MSE and MAE metrics, age estimation typically prioritizes the MAE metric. We augmented L1, Huber, and L2 loss functions with both original pairwise learning (L2+PRL) and AdaPRL to assess performance across two datasets. The experimental findings yielded the following insights: within the AgeDB and CLAP2016 datasets, AdaPRL consistently attained the most exemplary performance irrespective of the baseline regression loss employed. Specifically, on the AgeDB dataset, the MAE metric was reduced by 1.13%, 0.36%, and 0.71% when combined with the L1, Huber, and L2 regressions, respectively; on the CLAP2016 dataset, the reductions were 7.47%, 6.13%, and 4.13%, respectively. Additionally, while achieving more accurate age estimations, both PRL and AdaPRL also gained improvements in the Kendall’s $\tau$ across both datasets. This demonstrated that our approach excelled in preserving the ordinal relationships of actual ages, thereby more precisely reflecting the relative age dynamics among individuals.
@@
Table 10: Comparison of various methods’ performance on the AgeDB and CLAP2016 datasets. Rel.Imp denotes the relative improvement over the corresponding metrics.
A.4 Natural Language Understanding
@@ A.4.2 Experiment Setup
In our experiments, we employed a pre-trained RoBERTa Liu et al. (2019) base model as the baseline backbone. To better facilitate training in larger batch sizes and improve accuracy, we used the LAMB You et al. (2020) optimizer with a batch size of 320 and a learning rate of 8e-4 using a warm-up cosine annealing scheduler. We implemented an early stopping strategy based on a patience of 2 in the validation set. A grid search was conducted on the hyperparameters including ranking loss type, $\alpha$ between {0.001,0.002, 0.005, 0.01,0.02,0.05,0.1,0.2} and $\theta$ between {0,0.3,0.9}. The experiments were carried out over 10 runs to reduce randomness, and the average performance on the metrics of the test set was reported.
@@ A.4.3 Overall Performance
The detailed experimental results are presented in Table 11. Similarly, we emphasized the top-performing datasets in bold and denoted the second-best performing datasets with underlines. The findings demonstrated that the proposed AdaPRL(L2) loss function excelled, achieved the lowest MSE of 0.3561 and MAE of 0.2922, and represented improvements of 3.39% and 2.92% over the standard L2 loss respectively. Although the original pairwise learning (L2+PRL) did not match AdaPRL’s performance, it still offered a significant enhancement over the L2 loss in both MSE and MAE. Additionally, we can also observe that both AdaPRL and original pairwise learning (L2+PRL) attained a higher Kendall’s $\tau$ (0.6858) compared to L2, further illustrating that our approach more effectively maintains the relative ranking among samples when predicting the toxicity levels of user comments, which is crucial for accurately assessing the severity of reviews.
@@
Table 11: Regression accuracies for Pairwise Regression Learning and AdaPRL on Jigsaw. Rel.Imp indicates relative improvement over corresponding metrics. The experiments were repeated 10 times, and the average results were reported.
@@ A.5.1 Datasets
Janestreet [^8] https://www.kaggle.com/competitions/jane-street-real-time-market-data-forecasting/data Desai et al. (2024) data comprise a set of time series with 79 anonymized features, 1 sample weight column, and 9 responders, representing real market data. The goal is to forecast one of the responders for the future. We formalize the task as a univariate regression problem, skip the first 500 days of data to avoid missing data, and randomly split 50% / 50% of the data as validation and test set for the last 100 days. Following the data description and the official evaluation protocol, we use weighted $R^{2}$ as evaluation metrics.
@@ A.5.2 Experiment Setup
In our experiments, we employed a multi-head attention model Vaswani et al. (2023) as encoder followed by a multilayer perceptron as the baseline backbone. The input sequence length, stride, and prediction length were set to 64, 16, and 1 respectively, and the model was trained on respond_6 as a single target. All experiments used AdamW Loshchilov (2017) as optimizer and the weighted MSE as $L_{reg}$. We implemented an early stopping strategy based on a patience of 3 in the validation set. A grid search was conducted on the hyperparameters including ranking loss type, $\alpha$ between {0.005, 0.01,0.02,0.04,0.08,0.16,0.32}. The experiments were conducted over 3 runs and we reported the average weighted $R^{2}$ metrics on the test set.
@@ A.5.3 Overall Performance
The comprehensive experimental results are delineated in Table 12. The premier method AdaPRL attained the highest weighted $R^{2}$ score of 0.00617, thereby eclipsing alternative approaches. In comparison to the baseline loss function weighted MSE, which yielded a weighted $R^{2}$ of 0.00540, AdaPRL demonstrated a relative enhancement of 14.81%. In addition, the pairwise learning technique weighted MSE+PRL achieved a marginally superior performance of 0.00575, yet it did not match the precision of AdaPRL. These findings substantiate that AdaPRL effectively increases regression accuracy, surpassing both the conventional weighted MSE and the standard pairwise learning-based approaches.
@@
Table 12: Regression accuracies for Pairwise Regression Learning (PRL) and AdaPRL on Janestreet. Rel.Imp indicates relative improvement over corresponding metrics. The experiments were repeated three times, and the average results were reported.
@@ A.6.2 Overall Performance
As shown in Table 13, AdaPRL achieved the highest accuracy among various loss functions on our industrial dataset, demonstrating the effectiveness of the proposed training techniques. Native pairwise regression learning improved MSE and MAE by 0.51% and 1.60%, respectively. In contrast, the AdaPRL method, leveraging pairwise learning and uncertainty estimation, achieved total gains of 0.96% and 1.93% in MSE and MAE. In particular, in large-scale music streaming applications with tens of millions of users, even a modest improvement in accuracy of 0.1% is empirically significant and can positively impact business outcomes. The proposed method showcases substantial superiority in capturing and utilizing the relative relationships within user data. Furthermore, Kendall’s $\tau$ score for the proposed approach is the highest among the evaluated loss functions, reflecting enhanced ranking capabilities, which are crucial for identifying high-value users in real-world applications.
@@
Table 13: Regression accuracies for Pairwise Regression Learning (PRL) and AdaPRL on large-scale proprietary industrial regression dataset from a major music streaming app. Rel.Imp indicates relative improvement over corresponding metrics. The experiments were repeated three times and the average results were reported.
@@ B.1 Sensitivity of Hyperparameter Alpha
B.1.1 Experiment Setup
To investigate the impact and sensitivity of the hyperparameter in AdaPRL, we conducted experiments on the hyperparameter alpha in this section. We chose the MovieLens-1M, KuaiRec, and ETT datasets for analysis. We employed an RMSE type AdaPRL as the auxiliary loss for the MovieLens-1M dataset and MAE type AdaPRL for all the other datasets. Note that $\theta$ was set to 0 for all experiments. All other experimental settings remained consistent with those described in A.
@@ B.1.1 Experiment Setup
B.1.2 Results
The results are shown in Figures 2 and 3. As illustrated in the figures, AdaPRL consistently achieved higher precision than the regression baseline on the MovieLens-1M, KuaiRec, and ETT datasets in most choices of alpha. Moreover, as shown in our earlier experiment sections, a coarse-grained search on alpha yielded a significant improvement with AdaPRL on multiple regression datasets. These results suggest that our proposed AdaPRL approach is robust and not highly sensitive to the choice of the alpha hyperparameter.
@@ B.2 Effect of Sparsity in AdaPRL
B.2.1 Experiment Setup
As we analyzed before in 2.3.2, if randomness is induced for $L_{SCPRL}$ when generating the sparse area $S$, this method could be viewed as a variant of dropout applied to confidence pairwise learning. Therefore, we found it essential to perform a quantitative analysis of the potential effects of dropout on the algorithm and empirically investigated how the underlying mechanisms of random dropout influenced the model. The evaluation was conducted on a subset of the ETT dataset, specifically focusing on prediction tasks with a forecasting horizon of 720 time steps in this experiment to analyze since the problem of the bottleneck of memory and computation was mainly raised in the scenario of multivariate time series forecasting while incorporating AdaPRL. We used iTransformer as the baseline model and conducted a grid search of sparsity based on random masking techniques; hyperparameter alpha was optimized via a coarse-grained grid search, and the best result was reported. All other settings remained the same as were described above in A.2.2.
@@ B.2.1 Experiment Setup
B.2.2 Results
The result is shown in Figure 4. It is obvious that as the sparsity of the area gradually decreases from 0.9 to 0.1, where computation gradually increases, the algorithm’s evaluation loss MAE and MSE exhibited an overall downward trend. This indicates that reducing sparsity improves the model’s predictive accuracy and robustness, likely by preserving essential information and mitigating the loss of critical features. However, when the sparsity is set to 1.0, corresponding to using a non-sparse algorithm, both MAE and MSE are worse than those observed at most other sparsity levels. This suggests that the introduction of a certain (probably relatively lower) level of sparsity, corresponding to a small drop rate in dropout, plays a crucial role in preventing overfitting and reducing computational complexity, thereby improving generalization performance. Conversely, the absence of sparsity, although it computes the relative difference of more pairs of data and requires more memory consumption, still leads to suboptimal performance compared with a sparse version of AdaPRL. The complete result of the sparsity test of the ETT dataset is listed in Table 14.
@@ B.3 Generalization and Integration with Existing Methods
B.3.1 Experiment Setup
In this section, we evaluated the generalizability of AdaPRL and its compatibility with previously proposed methods. As shown in Table 1, AdaPRL demonstrated a significant improvement over the traditional L2 loss function across the majority of regression datasets. However, when comparing AdaPRL employed directly as $L_{reg}$ with the RankSim and Ordinal Entropy methods, AdaPRL performed slightly inferiorly on certain datasets such as KuaiRec, AgeDB and Jigsaw. Consequently, it was essential to explore whether integrating RankSim and Ordinal Entropy with AdaPRL could further boost the performance. Therefore, we designed new experiments on these three datasets by combining AdaPRL with both RankSim and Ordinal Entropy, aiming to assess the resultant performance.
Specifically, the hyperparameters of both RankSim and Ordinal Entropy method were optimized in each dataset via a coarse-grained grid search, and the weights were searched across {0.05,0.1,0.2,0.4,0.8,1.6}. Secondly, as we described in algorithm1, we computed and added an alpha-weighted $L_{CPRL}$ to the loss function of RankSim and Ordinal Entropy with the corresponding optimized hyperparameter. Then a coarse-grained grid search was performed on $\alpha$ of AdaPRL. The grid search and other settings were kept the same as the earlier experiment details.
B.3.2 Results
The results are presented in Table 15. As shown in the table, the integration of AdaPRL with both RankSim and Ordinal Entropy exhibited performance improvement compared to those without AdaPRL in the KuaiRec, AgeDB as well as Jigsaw datasets. Specifically, for RankSim method, the improvement of combining AdaPRL in MSE and MAE metrics was 0.03%, 1.04% for KuaiRec, 0.77%, 0.17% and 0.09% for AgeDB, and 0.23% and 5.46% for Jigsaw. The improvement for Ordinal Entropy was 6.46% and 16.72% for KuaiRec, 1.15% and 0.66% for AgeDB, and 2.22% and 2.57% for Jigsaw, respectively. The results demonstrate that our proposed AdaPRL is not a substitute for these existing techniques. Instead, it can be seamlessly integrated into various regression frameworks to achieve further performance improvements.